OpenStudy (anonymous):

What number should be added to each of the three numbers 1, 7, and 19 so that the resulting three numbers form a geometric progression?

OpenStudy (anonymous):

The answer is 5 but i'm not sure how to get there.

OpenStudy (anonymous):

If the resulting numbers (a1, a2 and a3) form a geometric progression they must fulfill: a2=a1·r--->a2/a1=r a3=a2·r--->a3/a2=r We can say: a2/a1=a3/a2 <--->a2·a2=a1·a3 (1) But we know that a1=1+x a2=7+x a3=19+x Then we can replace in (1) (7+x)(7+x)=(1+x)(19+x)---> 49+14x+x^2=19+20x+x^2--->49+14x=19+20x---> 49-19=30=20x-14x=6x--->6x=30--->x=5

OpenStudy (anonymous):

easy, isnt it?

OpenStudy (anonymous):

Much easier when we simplify it to a1 a2 and a3 before approaching it! Thank you so much!

OpenStudy (anonymous):

you are welcome